### v1

# Efficient Fine-Tuning of Quantized Models via Adaptive Rank and Bitwidth

Abstract
QLoRA effectively combines low-bit quantization and LoRA to achieve memory-friendly fine-tuning for large language models (LLM). Recently, methods based on SVD for continuous update iterations to initialize LoRA matrices to accommodate quantization errors have generally failed to consistently improve performance. Dynamic mixed precision is a natural idea for continuously improving the fine-tuning performance of quantized models, but previous methods often optimize low-rank subspaces or quantization components separately, without considering their synergy. To address this, we propose QR-Adaptor, a unified, gradient-free strategy that uses partial calibration data to jointly search the quantization components and the rank of low-rank spaces for each layer, thereby continuously improving model performance. QR-Adaptor does not minimize quantization error but treats precision and rank allocation as a discrete optimization problem guided by actual downstream performance and memory usage. Compared to state-of-the-art (SOTA) quantized LoRA fine-tuning methods, our approach achieves a 4.89% accuracy improvement on GSM8K, and in some cases even outperforms the 16-bit fine-tuned model while maintaining the memory footprint of the 4-bit setting.
1 Introduction
Large Language Models (LLMs) have achieved remarkable success in both language understanding and generation [Makridakis et al., 2023, Raiaan et al., 2024, Chang et al., 2024], yet adapting these powerful models to specific downstream tasks is often hindered by immense computational and memory costs [Wan et al., 2023]. Parameter-Efficient Fine-Tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) [Hu et al., 2022], address these bottlenecks by introducing lightweight low-rank updates to the core model parameters, thereby reducing the storage and compute needed for full gradient-based optimization. In parallel, quantization techniques [Gong et al., 2014, Gupta et al., 2015] compress weight matrices to fewer bits, enabling faster inference and lower memory footprints. Building on these two lines of research, QLoRA [Dettmers et al., 2023] has become a widely adopted paradigm for memory-efficient LLM fine-tuning, effectively pairing quantization with LoRA.
Recent works aim to further reduce quantization errors by adjusting the quantized weights before fine-tuning. For instance, LoftQ [Li et al., 2023] fits a portion of the error $\mathbf{W}-\mathbf{Q}$ through SVD-based initialization, and LQ-LoRA [Guo et al., 2024] attempts a similar correction strategy. While these approaches can boost performance after one iteration, multiple iterative corrections sometimes degrade accuracy—as illustrated in Figure 1, where LoftQ performance declines over several rounds. This drop arises in part from a mismatch between the discrete nature of quantization and continuous SVD updates, which risk shifting the model away from both $\mathbf{Q}$ and $\mathbf{W}$. Additionally, methods that dynamically adjust LoRA’s rank at run time, such as AdaLoRA [Zhang et al., 2023b], may introduce extra complexity for quantized models whose robustness has already been weakened [Gong et al., 2024].
Figure 1: LoftQ iterations, quantization error, fine-tuning performance, and time cost. As the iteration count increases, the overall quantization error slightly decreases but does not yield proportional boosts in downstream performance. Moreover, the fraction of the error that can still be reduced shrinks with each iteration, ultimately fitting around 20% of the quantization error.
We introduce QR-Adaptor, an approach that aims to circumvent these shortcomings by exploring each layer’s bitwidth and LoRA level in a uniform search space. Rather than optimizing quantization and LoRA parameters in isolation—or iteratively refining continuous residuals—QR-Adaptor jointly determines each layer’s precision level ($q_{l}$) and adaptation capacity ($r_{l}$). This approach exploits the interplay between numerical precision and low-rank expressiveness: certain layers may demand a higher bit-width to preserve critical features, while others benefit more from an increased LoRA rank to capture complex task-specific cues. By considering both factors together, QR-Adaptor achieves more robust control over memory usage and downstream accuracy. Specifically, our method conducts a gradient-free, multi-objective search over discrete $(q_{l},r_{l})$ pairs for each layer. To keep the computational burden comparable to a single SVD iteration, we evaluate candidate configurations only on a small partial calibration subset of the training data. Through this design, QR-Adaptor avoids the shortcomings of iterative SVD-based error fitting, and leverages a cohesive search strategy that systematically aligns quantization precision with LoRA’s representational power in quantized LLM fine-tuning.
To robustly explore the high-dimensional configuration space, we adopt two established multi-objective optimization methods. First, we perform a global exploration using a Pareto-Ranking Genetic Algorithm reminiscent of NSGA-II Deb et al. [2002], which maintains a diverse population of candidate solutions. Second, we apply local refinement with Bayesian optimization based on Gaussian processes, homing in on locally optimal trade-offs between performance and memory. By carefully integrating these techniques, QR-Adaptor systematically avoids the shortcomings of iterative error-fitting, improving upon baseline approaches in both efficacy and versatility. We summarize our main contributions as follows:
  • We uncover the reasons why existing iterative methods, which focus on reducing quantization errors through continuous optimization, fail to consistently enhance the performance of quantized models. In particular, we find that merely refining SVD-based corrections or individually adjusting bit-widths or LoRA ranks cannot adequately boost the capabilities of quantized LLMs.
  • We propose a multi-objective, gradient-free optimization strategy, QR-Adaptor, which jointly searches for the per-layer bit-width and LoRA rank. This approach integrates task-informed initialization, a Pareto-ranking genetic algorithm, and Bayesian optimization, and keeps the overhead comparable to a single SVD iteration by using calibration data.
  • Our method consistently outperforms previous baselines on multiple benchmarks (e.g., achieving a 4.89% improvement in accuracy on GSM8K). In some cases, it even surpasses 16-bit fine-tuned models while requiring memory comparable to that of a 4-bit setting.
2 Background
In this section, we analyze why iterative methods that aim to reduce quantization errors via continuous SVD-based updates can be suboptimal when weights are constrained to a discrete lattice. We also provide optimization paths for three different fine-tuned initialization strategies in Appendix 7, illustrating that iterative fitting of quantization errors may not be able to approximate the original accuracy model, thus motivating our proposed approach.
2.1 Discrete Nature of Quantization
Definition 2.1 (Quantization Operator).
For a weight matrix $\mathbf{W}\in\mathbb{R}^{d\times k}$ and quantization step $\Delta$, the quantization operator is defined as:
$\text{Quantize}(\mathbf{W})\;=\;\Delta\cdot\left\lfloor\frac{\mathbf{W}}{\Delta}+\tfrac{1}{2}\right\rfloor,$ (1)
[…]
$\mathcal{L}\;=\;\Delta\cdot\mathbb{Z}^{d\times k}\;\subset\;\mathbb{R}^{d\times k}.$ (2)
Interpretation.
This formalizes that each parameter in $\mathbf{W}$ can only take values from a discrete set (the “lattice”) after quantization. Even small changes in real-valued space may map to the same discrete point under rounding, causing discontinuities in any gradient-based updates.
2.2 Mismatch in Continuous Approximations
Existing methods often minimize
$\min_{\mathbf{A,B}}\Bigl{\|}\mathbf{W}-\underbrace{\text{Quantize}(\mathbf{W})}_{\in\mathcal{L}}\;-\;\mathbf{A}\mathbf{B}\Bigr{\|}_{F}.$ (3)
However, this objective presumes a continuous measure of error (the Frobenius norm) even though the feasible set $\text{Quantize}(\mathbf{W})$ is discrete.
Key Issues.
  • Space Duality: The solution space for $\mathbf{A}\mathbf{B}$ is continuous ($\mathbb{R}^{d\times r}\times\mathbb{R}^{r\times k}$), whereas $\text{Quantize}(\mathbf{W})$ must remain in the discrete lattice $\mathcal{L}$. Thus, $\text{Quantize}(\mathbf{W})+\mathbf{AB}$ might lie outside $\mathcal{L}$ after updates, negating any continuous gains.
  • Gradient Discontinuity: Let $\mathcal{P}_{\mathcal{L}}$ project onto $\mathcal{L}$. True gradient flow would require differentiating through $\mathcal{P}_{\mathcal{L}}(\text{Quantize}(\mathbf{W})+\mathbf{AB})$. But $\mathcal{P}_{\mathcal{L}}$ has zero gradient almost everywhere, making standard backpropagation ineffective for adjusting $\mathbf{AB}$ in a discrete context.
Implication.
Because $\mathbf{Q}=\text{Quantize}(\mathbf{W})$ is fixed to a discrete set, continuously optimizing $\mathbf{A,B}$ to reduce $\|\mathbf{W}-\mathbf{Q}-\mathbf{A}\mathbf{B}\|_{F}$ can yield diminishing or no improvements in final performance. Each step might look beneficial in a continuous sense but fails to move to a truly better discrete state.
2.3 Formal Proof of Suboptimality
Proposition 2.2.
Let
$\mathbf{AB}^{*}\;=\;\operatorname*{arg\,min}_{\mathbf{A,B}}\Bigl{\|}\mathbf{W}\;-\;\text{Quantize}(\mathbf{W})\;-\;\mathbf{A}\mathbf{B}\Bigr{\|}_{F}.$ (4)
Then there exists a $\mathbf{Q}^{\prime}\in\mathcal{L}$ such that
$\|\mathbf{W}-\mathbf{Q}^{\prime}\|_{F}\;<\;\|\mathbf{W}-\text{Quantize}(\mathbf{W})-\mathbf{AB}^{*}\|_{F}.$ (5)
Proof.
By the projection theorem:
$\|\mathbf{W}-\mathcal{P}_{\mathcal{L}}(\mathbf{W})\|_{F}$ $\;\leq\;\|\mathbf{W}-\mathbf{X}\|_{F}\quad\forall\mathbf{X}\in\mathcal{L},$ (6)
$\|\mathbf{W}-\mathbf{Q}^{\prime}\|_{F}$ $\;\leq\;\|\mathbf{W}-\text{Quantize}(\mathbf{W})-\mathbf{AB}^{*}\|_{F},$ (7)
where $\mathbf{Q}^{\prime}=\mathcal{P}_{\mathcal{L}}(\mathbf{W})$. Equality holds only if $\text{Quantize}(\mathbf{W})+\mathbf{AB}^{*}\in\mathcal{L}$, requiring $\mathbf{AB}^{*}$ to exactly compensate discrete rounding errors—an event with measure zero in high dimensions. ∎
Interpretation.
This proposition states that even the best continuous low-rank correction $\mathbf{AB}^{*}$ cannot fully offset the errors introduced by discrete quantization. A strictly better discrete point $\mathbf{Q}^{\prime}$ often exists in lattice space, demonstrating that methods trying to fix $\text{Quantize}(\mathbf{W})$ plus a real-valued $\mathbf{A}\mathbf{B}$ can remain suboptimal. In simpler terms, no matter how well we approximate $\mathbf{W}-\text{Quantize}(\mathbf{W})$ by a continuous matrix, there may be a better purely discrete solution.
2.4 Manifold Analysis
We can decompose the parameter space as
=$\mathcal{M}\;=\;\mathcal{L}\;\oplus\;\bigl{\{}\mathbf{AB}\,\mid\,\text{rank}(\mathbf{AB})\leq r\bigr{\}}.$ (8)
Because $\mathcal{L}$ is discrete and $\{\mathbf{AB}\}$ is continuous, their sum $\mathcal{M}$ is a hybrid space.
Critical Observations.
  • The intersection $\mathcal{L}\,\cap\,(\text{Quantize}(\mathbf{W})+\{\mathbf{AB}\})$ forms isolated points, underscoring the rare alignment needed for continuous $\mathbf{AB}$ to fix discrete rounding completely.
  • The effective search space $\mathcal{M}_{\text{eff}}=\{\mathcal{P}_{\mathcal{L}}(\text{Quantize}(\mathbf{W})+\mathbf{AB})\}$ has a fractal-like structure, reflecting the discontinuous jumps among discrete levels.
Implication.
As dimension $d\times k$ grows, the measure of $\mathbf{AB}$ that perfectly lands in the same discrete cell as $\mathbf{W}$ becomes negligible. This reinforces the statement that purely continuous approximations rarely coincide with actual quantized solutions, making Frobenius-based SVD updates unreliable for improving final performance.
2.5 Dynamical System Perspective
Training dynamics can be modeled by
$\frac{d}{dt}\begin{pmatrix}\mathbf{A}\\[4.0pt] \mathbf{B}\end{pmatrix}\;=\;-\eta\,\nabla_{\mathbf{A,B}}\,\mathcal{L}\Bigl{(}\mathcal{P}_{\mathcal{L}}\bigl{(}\text{Quantize}(\mathbf{W})+\mathbf{AB}\bigr{)}\Bigr{)}.$ (9)
Key Properties.
  • The right-hand side is piecewise constant, with jumps on the boundaries of $\mathcal{L}$. Small changes in $\mathbf{A},\mathbf{B}$ that do not cross a boundary in $\mathcal{L}$ produce no difference in the discrete projection.
  • Eigenvalues of the Jacobian vanish when projected onto the interior of a lattice cell. This means typical continuous gradient-based methods cannot easily escape suboptimal cells since they see "flat" gradients within each cell.
Meaning for Practical Fine-Tuning.
This piecewise constant nature implies that iterative SVD or gradient-based fitting will exhibit plateau-like behavior in large regions, only changing when crossing quantization boundaries. Consequently, even if $\|\mathbf{W}-(\mathbf{Q}+\mathbf{AB})\|_{F}$ looks partially improved in a real-valued sense, the actual quantized weights remain stuck. Hence, iterative error-fitting procedures may end up stagnating.
Collectively, these findings explain why repeated SVD-based steps may reduce the continuous norm $\|\mathbf{W}-\mathbf{Q}\|_{F}$, yet fail to yield stable accuracy improvements under quantization. In contrast, our approach (§3) avoids this pitfall by adopting a gradient-free, multi-objective strategy. Instead of directly minimizing Frobenius error in a continuous space, we evaluate each candidate (bit-width, LoRA rank) within the discrete quantized domain to optimize real downstream performance and memory usage.
3 Methodology
3.1 Problem Formulation
We consider a pre-trained LLM of $L$ layers. Each layer $l$ has a base weight matrix $\mathbf{W}_{l}\in\mathbb{R}^{d\times k}$, which we wish to both quantize to a lower bit-width $q_{l}$ and augment with a low-rank LoRA adaptation of rank $r_{l}$. Specifically, let
$q_{l}\in\mathcal{Q}\quad\text{and}\quad r_{l}\in\mathcal{R},$
where $\mathcal{Q}$ is the set of all feasible bit-width values (e.g., 4, 8 bits) and $\mathcal{R}$ the set of all feasible LoRA ranks. We unify these layer-wise assignments into a single configuration:
$C\;=\;\bigl{\{}(q_{1},\,r_{1}),\,(q_{2},\,r_{2}),\;\dots,\;(q_{L},\,r_{L})\bigr{\}},$
and let $\mathcal{C}$ be the space of all such configurations.
Layer-wise Quantization & LoRA.
For layer $l$, we apply a quantization function to obtain:
$\hat{\mathbf{W}}_{l}^{\,q_{l}}\;=\;\text{Quantize}\bigl{(}\mathbf{W}_{l},\,q_{l}\bigr{)}.$ (10)
We then add a low-rank term $\Delta\mathbf{W}_{l}^{\,r_{l}}=\mathbf{A}_{l}\,\mathbf{B}_{l}$, where $\mathbf{A}_{l}\in\mathbb{R}^{d\times r_{l}}$ and $\mathbf{B}_{l}\in\mathbb{R}^{r_{l}\times k}$. The layer’s forward pass for an input $\mathbf{x}\in\mathbb{R}^{k}$ is:
$\mathbf{y}\;=\;\hat{\mathbf{W}}_{l}^{\,q_{l}}\,\mathbf{x}\;+\;\Delta\mathbf{W}_{l}^{\,r_{l}}\,\mathbf{x}.$ (11)
Hence, each layer’s parameters become a combination of discrete weights $\hat{\mathbf{W}}_{l}^{\,q_{l}}$ (bit-width $q_{l}$) and a continuous LoRA update of rank $r_{l}$.
Objective: Performance vs. Memory.
Our goal is to find a configuration $C\in\mathcal{C}$ that yields high accuracy on a training set $\mathcal{D}_{\text{train}}$ (evaluated on downstream tasks $\mathcal{D}_{\text{test}}$) while minimizing memory usage. Let $P(C)$ be the model’s performance under configuration $C$, and let $M(C)$ be its total memory footprint (including quantized weights and LoRA parameters). We combine these criteria in a single weighted objective:
$\max_{C\,\in\,\mathcal{C}}\quad\alpha\,\frac{\,P(C)\,-\,\mu_{P}\,}{\sigma_{P}}\;-\;(1-\alpha)\,\frac{\,M(C)\,-\,\mu_{M}\,}{\sigma_{M}},$ (12)
$\text{subject to}\quad q_{l}\in\mathcal{Q},\;r_{l}\in\mathcal{R},\;\text{for each}\;l=1,\dots,L.$
Here, $(\mu_{P},\sigma_{P})$ and $(\mu_{M},\sigma_{M})$ are the mean and standard deviation of performance and memory, computed over sample configurations to normalize their scales. The parameter $\alpha\in[0,1]$ governs the trade-off: a larger $\alpha$ places more emphasis on performance $P(C)$, whereas a smaller $\alpha$ prioritizes memory reduction $M(C)$.
3.2 QR-Adaptor Framework
Our approach differs from previous methods relying on fixed or hierarchical dynamic single configurations. We jointly explore the configuration space of quantization bits and LoRA ranks, creating a comprehensive search space that encompasses all potential optimal configurations. The main challenges in implementing this gradient-free optimization process are (a) The high-dimensional, discrete nature of the configuration space. (b) The computational cost of evaluating performance. To address these challenges, we propose QR-Adaptor, a method that effectively finds the relative optimal solution in three stages.
Task Information Based Initialization
Our optimization pipeline starts by estimating the relative importance of each layer to the downstream task. Rather than relying on gradient norms (which may fail to capture a layer’s true influence on inference), we adopt an information-entropy criterion:
$I(l)\;=\;H(Y)\;-\;H\bigl{(}Y\mid X_{l}\bigr{)},$ (13)
where $H(Y)$ is the entropy of the model’s outputs, and $H(Y\mid X_{l})$ is the conditional entropy given the $l$-th layer’s intermediate representation. Layers with a higher $I(l)$ exert greater impact on final predictions. Empirically, we find (Appendix 13.1) that this entropy-based approach outperforms gradient-norm heuristics in selecting effective layers for bit-width and rank adjustments.
Concretely, we initialize per-layer quantization bits $\{q_{l}\}$ by assigning larger $q_{l}$ to layers with higher importance scores and smaller $q_{l}$ to those deemed less critical. After quantizing the model with these initial bits, we recompute layer importance under quantization and then allocate LoRA ranks $\{r_{l}\}$: layers that remain most vital post-quantization receive larger $r_{l}$, while others get reduced ranks. This two-step procedure narrows the search space to promising configurations from the outset, guiding subsequent multi-objective optimization toward higher accuracy and lower memory usage.
Figure 2: Detailed PRGA flow chart. The input is a set of solutions from the initialization, and the output is a set of Pareto front solutions containing multiple solutions.
Global Exploration with PRGA
In LLM fine-tuning, quantization bit and LoRA rank can be conceptualized as genes, with the resulting performance and memory usage analogous to phenotypic expressions in a population. Inspired by NSGA-II’s Deb et al. [2002] proven success in multi-objective optimization, we adapted its mechanisms to develop PRGA (Pareto Ranking Genetic Algorithm). PRGA explores the combined solution space of bits and ranks to identify the optimal Pareto frontier, simultaneously balancing performance and memory usage in LLM fine-tuning. This efficient multi-objective optimization algorithm uses an elitist selection approach to evolve a population of configurations, each represented as $C=\{(q_{l},r_{l})_{l=1}^{L}\}$, where $q_{l}$ and $r_{l}$ denote the quantization bits and LoRA rank for layer $l$, respectively. PRGA iteratively applies selection, crossover, and mutation operations to the population, aiming to simultaneously maximize performance and minimize memory usage. The algorithm progresses until it reaches a predefined stopping criterion or a maximum number of iterations, whichever comes first. Once finished, the algorithm ultimately produces a Pareto frontier that represents the optimal trade-offs between our competing objectives.
Before introducing the PRGA flow, we first introduce some key foundational concepts. In a multi-objective minimization problem with $n$ objective components $f_{i}(x),i=1,\ldots,n$, the Pareto Dominance Relationship is defined between any two decision variables $X_{a}$ and $X_{b}$. We say that $X_{a}$ dominates $X_{b}$ if for all $i\in\{1,2,\ldots,n\}$, $f_{i}(X_{a})\leq f_{i}(X_{b})$, and there exists at least one $i\in\{1,2,\ldots,n\}$ such that $f_{i}(X_{a})<f_{i}(X_{b})$. A Non-dominated Solution is a decision variable that is not dominated by any other decision variable in the set. The concept of Pareto Rank is used to categorize solutions within a set. Non-dominated solutions are assigned a Pareto rank of 1. After removing these rank 1 solutions from the set, the remaining non-dominated solutions are assigned a Pareto rank of 2. This process continues iteratively, assigning increasing ranks to subsequent layers of non-dominated solutions until all solutions in the set have been ranked.
As illustrated in Algorithm 7, we employ the Pareto Ranking to sort all individuals within the population. To address solutions with identical Pareto ranks, we use the crowding distance $d$ for further differentiation within each Pareto rank. The detailed calculation of the crowding distance is presented in Algorithm 7.
Elite retention, a method simulating natural elimination, is performed after calculating the Pareto rank and crowding distance of all individuals in a generation. This process begins by combining the parent and offspring populations into a merged population. To generate the next generation, we start with the lowest Pareto rank and transfer entire layers of individuals from the merged population to the new population, moving progressively to higher ranks. This continues until we reach a layer that cannot be fully accommodated in the new population. For this partially accommodated layer, we sort its individuals based on their crowding distance in descending order and add them sequentially to the new population until it reaches its full capacity.
For the crossover and mutation operations, we employ methods analogous to Simulated Binary Crossover (SBX) and Polynomial Mutation, respectively. These methods are adapted to operate on $L$ pairs of positive integers $(q_{l},r_{l})$. The detailed procedures for these operations are presented in Algorithm 9 for the crossover and Algorithm 9 for the mutation. By applying these adapted SBX and polynomial mutation operations, we can effectively evolve the population of solutions represented by integer pairs, balancing exploration and exploitation.
The PRGA process, shown in Figure 2, begins by generating an initial population of size $N$ through controlled random variations based on the previous stage’s configuration. It then creates offspring using selection, crossover, and mutation operations. The parent and offspring populations are combined and subsequently fastly fine-tuned and validated on a subset of the training dataset. The resulting performance and memory metrics are used to calculate Pareto ranks for each individual. Next, the algorithm applies an elite retention strategy combined with crowding distance calculation to select individuals for the new population. This cycle repeats, generating consecutive generations until the termination condition is met, effectively exploring the solution space to optimize both performance and memory usage simultaneously.
Local Refinement with Bayesian Optimization
Figure 3: Detailed Bayesian optimization flow chart. The input is the Pareto front solution set from the global search, and the output is a set of optimal solutions obtained according to the requirements.
While PRGA effectively explores the global configuration space, it may not precisely capture local optima near the Pareto front. To further refine these solutions, we employ Bayesian optimization, a technique renowned for its ability to optimize expensive black-box functions with uncertainty quantification. We initiate this process by utilizing the solutions from the PRGA-generated Pareto front as our initial sampling points. For each point, we use these configurations to quickly fine-tune the model and test to obtain actual performance and memory usage. We then compute its corresponding weighted objective function value $y$ using the predefined objective function (Equation 12) and specified weight preferences. These $y$ values serve a dual purpose: firstly, they are combined with the covariance matrix $K$, which is constructed using the radial basis function (RBF) kernel to quantify similarities between sample points, to build a Gaussian process model; secondly, they enable us to identify the best-performing point, which becomes the focal point for subsequent searches.
The next phase involves employing a random search strategy to select new sampling points in the vicinity of this top-performing configuration. For each newly selected sampling point $x^{*}$, we leverage the Gaussian process to estimate its predicted value and associated uncertainty using the following equations:
$\small\begin{split}\mu(x^{*})&=m(x^{*})+K(x^{*},X)K(X,X)^{-1}(y-m(X))\\ \sigma^{2}(x^{*})&=k(x^{*},x^{*})-K(x^{*},X)K(X,X)^{-1}K(X,xj)\end{split}$ (14)
where $m(x^{*})$ is the prior mean function, $K(x^{*},X)$ is the covariance between the new point and the existing points, and $K(X,X)$ is the covariance matrix of the existing points.
We then employ the Expected Improvement (EI) as the acquisition function, calculating it using the following formula:
$EI(x^{*})=\sigma(x^{*})\left(Z\cdot\Phi(Z)+\phi(Z)\right)$ (15)
$\text{and}\quad Z=\frac{\mu(x^{*})-y_{\text{best}}}{\sigma(x^{*})}.$
where $\Phi(Z)$ is the cumulative distribution function of the standard normal distribution, while $\phi(Z)$ is its probability density function and $y_{best}$ is the best objective value among all points.
We select the point with the largest EI value as the next evaluation point and quickly fine-tune the LLM using this configuration. We then test to obtain the performance and memory usage, which are used to calculate the objective function value, and this new point is compared with the previous best point. If it proves to be a better choice, it can be updated as the optimal point and the search for the next iteration proceeds near it; otherwise, we continue searching near the original optimal point. Regardless of whether the best point is updated, the new point and its objective function value are incorporated into the training dataset to update the Gaussian process model before beginning the next iteration. This process repeats until a predefined termination condition is met, such as reaching maximum iterations, meeting a convergence criterion, or hitting a time limit. The detailed flowchart is shown in Figure 3.
The solution set obtained through this Bayesian optimization process offers a refined representation of high-quality configurations, effectively capturing the trade-offs between performance and memory usage. This iterative refinement process culminates in a final set of configurations that represent the best balance of our objectives based on the specified preferences. By presenting these optimized configurations as the final output, our approach enables practitioners to directly choose suitable configuration that aligns with their specific performance requirements and memory constraints.
4 Evaluation
Table 1: Performance comparison of different methods across various bit-width configurations on Llama3.1-8B. Superscripts on LoftQ bits indicate the number of initialization iterations. Bold figures represent the best performance for a given model and task, while underlined figures indicate the second-best. Accuracy is reported as %.
Method Bit ARC(C) ARC(E) BoolQ GSM8K HellaS OBQA PIQA WinoG Average
Rank = 8 LoRA 16 56.14 83.88 83.18 54.36 79.44 45.20 82.10 75.30 69.95
QLoRA 8 57.08 83.46 82.48 53.75 79.63 46.00 82.10 74.59 69.89
QLoRA 4 54.35 82.41 82.08 44.35 78.82 44.20 81.50 73.64 67.67
AdaLoRA 16 52.90 81.99 81.87 50.57 78.65 45.00 81.34 73.95 68.28
AdaLoRA 8 52.90 81.86 82.05 49.96 78.65 44.80 81.34 74.43 68.25
AdaLoRA 4 51.28 80.98 80.61 37.83 77.36 42.80 80.74 72.53 65.51
LoftQ 41 54.86 82.74 82.26 51.40 78.65 46.00 81.45 73.24 68.82
LoftQ 45 52.65 81.82 81.53 39.65 78.50 43.40 81.39 72.69 66.45
LoftQ 410 51.88 81.31 79.66 38.44 78.01 43.20 81.12 71.98 65.70
QR-Adaptor 5.45 56.83 84.12 83.38 56.29 80.93 45.80 82.92 75.10 70.67
Rank = 16 LoRA 16 56.74 83.63 83.00 54.13 79.51 44.40 81.83 74.43 69.70
QLoRA 8 56.23 82.91 82.66 53.68 79.46 46.00 81.66 74.74 69.67
QLoRA 4 53.84 81.99 82.11 44.66 78.76 44.40 81.72 73.09 67.57
AdaLoRA 16 53.07 82.03 81.99 50.11 78.61 45.40 81.28 74.11 68.33
AdaLoRA 8 53.33 82.03 82.11 49.13 78.57 45.20 81.34 73.79 68.19
AdaLoRA 4 50.85 80.72 80.73 37.98 77.34 42.80 80.52 73.16 65.51
LoftQ 41 55.12 82.58 82.69 49.81 78.82 45.80 81.28 74.27 68.80
LoftQ 45 53.92 82.32 81.56 42.00 78.54 43.80 81.56 72.77 67.06
LoftQ 410 52.90 81.69 81.56 39.88 78.64 43.80 81.07 71.98 66.44
QR-Adaptor 5.45 56.83 84.12 83.38 56.29 80.93 45.80 82.92 75.10 70.67
In this section, we first introduce the experimental setup, including datasets, models, baselines, and implementation details. All hyperparameters aside from rank value and bit-width are kept consistent with the baselines. Additional results are provided in Appendix 10 due to space constraints:
  • Extended Experimental Results: Further results across different models are detailed in Appendix 10.1, Appendix 10.2, and Appendix 10.6 and Appendix 10.7.
  • Impact of Higher LoRA Ranks on Larger Datasets: Evaluations of our method on larger datasets with higher LoRA ranks are detailed in Appendix 10.3.
  • Training Time Comparison: We analyze the training time efficiency of our method in Appendix 10.4.
  • Fair Bit-width Configuration Comparison: A fairer comparison of quantization methods with consistent bit-width configurations is given in Appendix 10.5.
We conduct ablation studies to assess the impact of different components on performance. We demonstrate the necessity of the three-stage process by removing one stage at a time. We also perform a sensitivity analysis of PRGA hyperparameters, as presented in Appendix 13.
4.1 Experimental Setup
Datasets and LLMs.
We utilize the Alpaca52k and hc3 [Taori et al., 2023] for fine-tuning and evaluate the zero-shot performance of these LLMs on benchmarks including BoolQ [Clark et al., 2019], PIQA [Bisk et al., 2020], HellaSwag [Zellers et al., 2019], WinoGrande [Sakaguchi et al., 2021], ARC-easy [Clark et al., 2018], ARC-challenge [Clark et al., 2018], OpenbookQA [Mihaylov et al., 2018], and MMLU Hendrycks et al. [2021]. The models used in our experiments are Llama2 Touvron et al. [2023] and Llama3.1 Grattafiori et al. [2024].
Baselines.
We compare our method against several baselines: without tuning, LoRA Hu et al. [2022], QLoRA Dettmers et al. [2023], Adalora Zhang et al. [2023b], LoftQ Li et al. [2023], and LQ-LoRA Guo et al. [2024]. We evaluated the performance of LoftQ with different iteration numbers. For Adalora, which dynamically allocates ranks based on the average rank budget, we set the budget to 8 and 64. Finally, for LQ-LoRA, which allocates quantization bit-width based on the average weight bit-width budget and quantization error, we set the bit-width budget to 4.
Implementation Details.
We utilize the following configurations: PyTorch version 2.1.2, BitsandBytes library version 0.43.1, Transformers library version 4.41.0, PEFT library version 0.11.1, Optuna library version 3.6.1, CUDA version 12.4, GPU: NVIDIA L20 GPU. Operating System: Ubuntu. Concise implementation details are provided in the Appendix 12.We define the population size as 5 and generate 1 new offspring in each iteration. The second and third phases were iterated 5 times.
Figure 4: From left to right, the actual measured performance and memory usage of the configurations generated by QR-Adaptor, QR-Adaptor without stage1, QR-Adaptor without stage2, and QR-Adaptor without stage3 are shown. Different colors represent the configurations generated at different stages.
4.2 Main Results
We present the performance comparison of LLaMA3.1 on commonsense understanding tasks in Table 1. For practical deployment, we exclude the 2-bit configuration in mixed-precision setups, as it does not significantly reduce memory usage. Results with 2-bit configurations are shown in Appendix 10.7. Our method achieves or surpasses the performance of 16-bit fine-tuned models while maintaining the same memory footprint as the 4-bit models.
As mentioned, LoftQ outperforms 4-bit QLoRA after one iteration, but its performance degrades with more iterations. This is due to the mismatch between continuous error correction and discrete quantization. AdaLoRA, which dynamically adjusts LoRA rank during fine-tuning, is less effective for quantized models, as the quantization process introduces discrete constraints incompatible with dynamic rank changes. Additional results for longer training epochs are available in Appendix 10.6.
QR-Adaptor jointly optimizes bit-width and LoRA rank, balancing precision and adaptation capacity for superior performance. Our results confirm that jointly optimizing these parameters leads to better downstream performance. Furthermore, QR-Adaptor allocates resources more efficiently by assigning higher LoRA ranks to critical layers and higher precision to important layers, achieving high accuracy with low memory usage.
Due to page limitations, please see Appendix 10 for more experimental results and analysis.
4.3 Ablation Study
We use the WinoGrande benchmark as an example for the ablation study to evaluate the role of each stage in QR-Adapto. As shown in Figure 4, it is evident that excluding PRGA and Bayesian optimization leads to uneven exploration of the search space—one is too broad and the other too concentrated—since they represent the extrapolation and interpolation capabilities, respectively. Excluding stage 1 results in overly scattered exploration because PRGA starts from a random search without an initialization point. However, it still manages to explore the theoretically optimal region in the upper-left corner, demonstrating the strong capabilities of PRGA and Bayesian optimization. In contrast, the complete three-stage QR-Adaptor shows the advantage of first conducting a broad exploration around the initialization point, followed by interpolation near promising solutions to further optimize and identify better configuration.
5 Related Work
LLM Quantization. The field of LLM quantization has witnessed substantial progress, driven by the need for efficient model deployment. Recent research has introduced several innovative approaches. Frantar et al. [2023] have developed GPTQ, which achieves 4-bit precision with layer-wise quantization. Lin et al. [2023] have proposed AWQ, which improves accuracy for heavily quantized models. Yao et al. [2022] have introduced ZeroQuant, which preserves zero-shot capabilities at lower bit widths. Dettmers et al. [2022] have presented LLM.int8(), which enables 8-bit quantization for consumer hardware. Kim et al. [2023] have combined quantization with pruning and knowledge distillation in SqueezeLLM. Guan et al. [2024] have optimized the balance between compression and performance through mixed-precision quantization with APTQ. These developments significantly enhance the efficiency of LLMs.
Parameter Efficient Fine-Tuning. PEFT techniques have become crucial for enhancing LLMs without increasing inference overhead. Recent innovations have expanded the field. Dettmers et al. [2023] have introduced QLoRA, which combines 4-bit quantization with low-rank adapters. Li et al. [2023] have presented LoftQ, which alternates between quantization and low-rank approximation steps. Berman and Peherstorfer [2024] have introduced CoLoRA for accelerating the prediction of solution fields under new parameters. AdaLoRA Zhang et al. [2023a] proposes adaptive budget allocation for low-rank updates, while LQ-LoRA Guo et al. [2023] combines low-rank decomposition with quantization for efficient fine-tuning under memory constraints. Additionally, Zhou et al. [2024] have introduced RankAdaptor, which is a hierarchical dynamic low-rank adaptation method for structural pruned LLMs. These advancements offer innovative solutions for efficient LLM fine-tuning across diverse applications.
6 Conclusion
In this work, we propose QR-Adaptor, a unified, gradient-free method that uses partial calibration data to simultaneously optimize the precision and LoRA rank of each model layer. By focusing on the discrete nature of quantization and low-rank spaces and optimizing them within a task-driven framework, QR-Adaptor overcomes the limitations of iterative error-fitting techniques and rank-adaptive methods unsuitable for quantization. Our extensive experiments demonstrate that QR-Adaptor consistently outperforms existing baselines, achieving better performance than 16-bit fine-tuned models while maintaining a 4-bit memory footprint. These results highlight the importance of integrating quantization and low-rank matrices into a single, cohesive optimization process, driven by actual performance and memory efficiency.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.